For this quiz, you need the financial statements, found in the annual report, for Walgreens year 2009. All the ratios and FCF you calculate should be based on the financial statements for Walgreens, year 2009. Also, round to the second decimal your answers. If after rounding to the second decimal the ratios for Wal-Mart and Walgreens are equal, then, you can say there is no difference between them, if they differ even in the second decimal, there is a difference and one company is performing better. For questions 10 to 19 please refer to the following table with numbers for Wal-Mart. Only FCF shareholders and FCF all are in million dollars, the other numbers are ratios hence with no dimension. The annual report for Walgreens, year 2009, is in the folder 'Accounting and other materials WalMart 11648 shareholders 13435 all t to book ratio OE 0.21 OA 0.08 0.14 OIC 0.88 urrent ratio uick ratio 0.26 t turnover ratio 2.48 ixed asset turnover ratio 4.37 0.06 ting profit margin Debt to equity ratio 1.50 to equity ratio 0.48 0.28 yout ratio
QUESTION Calculate the Free Cash Flow for all, the one available to shareholders and bondholders. QUESTION 2 Calculate the market value of the firm in million dollars, use a price of $33.84 per share. Remeber, do not use diluted shares, use basic if you use the average shares outstanding. QUESTION 3 Calculate book to market ratio, use the market value you calculated in question 2 QUESTION 4 Calculate Return on Equity QUESTION Calculate Return on Assets (ROA) QUESTION Calculate Asset Turnover Ratio.
Answer #1

I have used Walgreens 2009 financing statement for the year ending 31 August 2009. All the below numbers are for FY19 (ending Aug-19)

1) Free cash flow for all :

FCF to all = Cash flow from operating + Cash flow from investing activities

= 3,744 - 1,274

FCF to all = $2,470 millions

2) Market capitalization at $33.84 share price

Average number of shares outstanding at of 31 August 2019 = 990 million

Market cap = 990 * 33.84

Market cap as of 31 August 2019= $33.5 billion

3) Book to market value ratio = Market cap / Book value of equity

= 33.5 / 14.4

Book to market value ratio = 2.33

4) Return on Equity = Net income / Book value of equity

= 2,006 / 14,376

Return on Equity = 13.95%

5) Return on Assets = Net income / Average total assets = Net income / (total assets in 2009 + total assets in 2008)/2

= 2,006 / (25,142+ 22,410) /2

Return on Assets = 8.44%

6) Asset turnover ratio = Sales in 2009 / Average assets

= 63,335 / (25,142+ 22,410) /2

Asset turnover ratio = 2.66 times